Dish N’ Dat Company Product-Line Income Statement For the Month Ended March 31, 2016 1 Bowls Plates Cups 2 Sales $70,800.00 $106,170.00 $31,420.00 3 Cost of goods sold 32,650.00 42,340.00 16,780.00 4 Gross profit $38,150.00 $63,830.00 $14,640.00 5 Selling and administrative expenses 27,790.00 42,350.00 16,560.00 6 Income from operations $10,360.00 $21,480.00 $(1,920.00) Fixed costs are 10% of the cost of goods sold and 50% of the selling and administrative expenses. Dish N’ Dat assumes that fixed costs would not be materially affected if the Cups line were discontinued. Required: A. Prepare a differential analysis dated March 31, 2016, to determine if Cups should be continued (Alternative 1) or discontinued (Alternative 2). Refer to the lists of Labels and Amount Descriptions for the exact wording of the answer choices for text entries. For those boxes in which you must enter subtracted or negative numbers use a minus sign. If there is no amount or an amount is zero, enter "0". A colon (:) will automatically appear if required. B. Should the Cups line be retained? Explain.
Answer #1
A
Continue Cups Discontinue Cups Differential Effect on Income
(Alternative 1) (Alternative 2) (Alternative 2)
Revenues 31420 0 -31420
Costs:
Variable cost of goods sold -15102 0 15102
Variable selling and administrative expenses -8280 0 8280
Fixed costs -9958 -9958 0
Income(loss) -1920 -9958 -8038
B
Cups line should be retained as discontinuing cup line will reduce income by $8038
Workings:
Variable cost of goods sold =16780*(1-10%)
Variable selling and administrative expenses =16560*(1-50%)
Fixed costs =(16780*10%)+(16560*50%)